OpenStudy (anonymous):

the perimeter of a rectangle equals the sum of the lengths of its four side. if the width of rectangle is 9 inches and its perimeter is 54 inches, find the length of the rectangle.

OpenStudy (anonymous):

let the length be x 9+9+x+x=54 2x=54-18 2x=36 x=18